# Qubit-assisted quantum metrology

### v1

@@ Abstract
We propose a quantum metrology protocol based on a two-step joint evolution of the probe system and an ancillary qubit and a single-shot projective measurement. With an optimized initialization of the ancillary qubit, the quantum Fisher information (QFI) about the phase parameter encoded in the probe system is found to be determined by the expectation value of the square of a time-optimized phase generator, independent of the probe state. Therefore, QFI can approach the Heisenberg scaling $N^{2}$ with respect to the quantum number $N$, even when the probe system is prepared in a classical state. We find that this scaling behavior is robust against the imperfections in preparing the ancillary qubit and controlling the evolution time. Using the time-reversal strategy, the classical Fisher information (CFI) in our metrology protocol is saturated with its quantum counterpart. Our work thus paves an economical way to realize the Heisenberg-scaling limit in metrology precision with no use of entanglement or squeezing.
@@ I Introduction
Quantum metrology aims to use various quantum resources including entanglement and squeezing to excess the standard quantum limit (SQL), where the uncertainty decreases with $1/\sqrt{N}$, on performing measurements of physical parameters Sun et al. (2010); Ma et al. (2011); Genoni et al. (2012); Escher et al. (2012); Zhong et al. (2013). Examples of these parameters are the phases of evolving quantum systems encoded in multiple applications Paris (2009); Helstrom (1969); Yurke et al. (1986), such as gravitational wave detection Caves (1981), biological sensing Taylor and Bowen (2016); Mauranyapin et al. (2017), atomic clocks Ludlow et al. (2015); Katori (2011), and magnetometry Jones et al. (2009). A conventional protocol for precision metrology reaching the Heisenberg limit (HL) Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020), where the uncertainty decreases with $1/N$, consists of a well-prepared probe system (in e.g., the NOON state), a definite unitary transformation to encode the phase parameter $\theta$ into the probe state, a practical measurement, and a suitable data processing to produce an estimator $\tilde{\theta}(x)$ about the parameter $\theta$ from the outcome $x$. The estimation precision is quantified by the standard deviation $\delta\theta=\langle[\tilde{\theta}(x)-\theta]^{2}\rangle^{1/2}$, which is under the constraint of the Cramér-Rao inequality Helstrom (1969); Holevo (1982): $\delta\theta\geq 1/\sqrt{F_{c}}$ with $F_{c}$ the classical Fisher information (CFI). Maximizing $F_{c}$ over all possible measurements gives rise to the quantum Fisher information (QFI) $F_{Q}$ and hence the quantum Cramér-Rao bound $\delta\theta_{\rm min}=1/\sqrt{F_{Q}}$ Braunstein and Caves (1994); Braunstein et al. (1996); Luo (2003); Pezzé and Smerzi (2009); Kacprowicz et al. (2010) on the attainable sensitivity to estimate $\theta$.
@@
The optimal precision $\delta\theta_{\rm min}$ generally improves with an increasing quantum number $N$ of the probe state employed in the measurement. With the probe system in separable states, we have $F_{Q}\sim N$ for the standard quantum limit. With particular entangled probe states, such as the NOON state, we have the so-called Heisenberg scaling $F_{Q}\sim N^{2}$ and hence the metrological precision attains HL Leibfried et al. (2004); Mitchell et al. (2004); Boto et al. (2000); Gerry (2000). However, the difficulty in obtaining NOON states with a large $N$ hinders its practical applications Bollinger et al. (1996); Walther et al. (2004); Afek et al. (2010); Zhang et al. (2013). An alternative method to improve the precision is to use the squeezed spin state Wineland et al. (1992); Kitagawa and Ueda (1993); Wineland et al. (1994); Goda et al. (2008), that is typically generated by the one-axis twisting (OAT) and two-axis twisting (TAT) interactions Liu et al. (2011); Zhang et al. (2017). The collective OAT interaction $H_{\rm OAT}\propto J_{z}^{2}$ gives rise to a squeezing degree $\propto 1/N^{2/3}$ for $N$ particles, which has been implemented in Bose-Einstein condensation (BEC) systems Gross et al. (2010); Riedel et al. (2010), trapped ions Bohnet et al. (2016); Lu et al. (2019), superconducting qubits Song et al. (2019); Xu et al. (2020), and warm vapors in glass cells Hosten et al. (2016); Bao et al. (2020); Duan et al. (2024). Under the TAT interaction $H_{\rm TAT}\propto J_{x}^{2}+J_{y}^{2}$, that is mainly reported in theoretical schemes Zhang et al. (2017); Borregaard et al. (2017), the squeezing degree $\propto 1/N$ approaches HL. It remains however a challenge in practice since the interaction Hamiltonian is rarely found in current platforms Helmerson and You (2001); Borregaard et al. (2017); Macrì et al. (2020). In quantum optics, the archetype of a metrology experiment for phase estimation is the Mach-Zehnder interferometer. When the beam splitters are replaced with optical parametric amplifiers, the resulting $SU(1,1)$ interferometer can reach HL only by inputting coherent states Yurke et al. (1986); Li et al. (2014, 2016); Jing et al. (2011); Hudelist et al. (2014); Du et al. (2022). However, its sensitivity to the internal losses eventually makes it work even below SQL Marino et al. (2012).
@@
Recently, the coupling with an ancillary system is found to outperform SQL in the estimation precision about the phase parameter. In a protocol for measuring the light rotation with indefinite time direction Xia et al. (2024), an ancillary system serves as a quantum switch leading to a superposition of opposite rotations of the probe system. The rotation precision can thus be enhanced to HL even by using classical probe states. And a precision of $12.9$ nrad on light rotation measurement has been experimentally realized with a $150$-order Laguerre-Gaussian beam. Motivated by the time reversal strategy Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024), based on which a quantum metrology protocol can approach the Cramér-Rao bound for certain parameters, Ref. Luo and Yu (2023) proposed a controlled protocol based on a two-photonic-mode system and an ancillary qubit. In this protocol, the interaction between the photonic system and the qubit can be manipulated by the external control fields to implement the unitary transformations along the forward- and reverse-time directions. The forward-time evolution generates a photonic entangled NOON state to improve QFI and the reverse-time evolution allows CFI to saturate with its quantum counterpart.
@@
In this work, we present a qubit-assisted metrology protocol to estimate the phase parameter $\theta$ encoded into the probe system (a spin ensemble) during a time evolution, which is based on the time-reversal strategy and a projective measurement. With a properly prepared ancillary qubit, QFI can be modified from the variance of the phase generator to the mean square of that with respect to the probe state. In this case, QFI can be enhanced up to the Heisenberg scaling in terms of the total spin number ($F_{Q}\propto N^{2}$) even when the probe spin ensemble is prepared as a classical state. The scaling behavior is robust against the imperfections about the initial state of the ancillary qubit and the optimized evolution time of the whole system. Moreover, CFI of our protocol is found to be coincident with QFI by the time-reversal strategy.
@@
The rest of this work is structured as follows. In Sec. II, we describe the circuit model of our metrology protocol with the ancillary qubit. In Sec. III, we investigate the conditions for the time-reversal strategy. A detailed derivation about the relevant evolution operator is provided in Appendix A. In Sec. IV, we find that QFI in our protocol can be expressed as the mean square of an optimized phase generator with respect to the probe initial state. That underlies the Heisenberg-limited sensitivity for a large $N$ even when the probe is in the mixed state. In Sec. V, we calculate CFI in our protocol as the extractable information from the probability distribution of the output state upon the projective measurement. In Sec. VI, we adapt our protocol to an alternative interaction between the probe system and an ancillary qubit. We summarize our work in Sec. VII.
II Hamiltonian and metrology model
@@
Consider a model consisting of a spin ensemble as a measurement probe and an ancillary spin-$1/2$. The full Hamiltonian including the free Hamiltonian $H_{0}$ and interaction part $H_{I}$ reads ($\hbar\equiv 1$)
@@
The full Hamiltonian (1) is feasible in various experiments. In quantum dots Gillard et al. (2022), $J_{z}$ and $\sigma_{z}$ describe the nuclear spins and the electron spin, respectively. In BEC, the Hamiltonian can be realized by two cavities containing two-component BECs coupled by an optical fiber Pyrkov and Byrnes (2013), where the operators $J_{z}$ and $\sigma_{z}$ correspond to the Schwinger boson operators $a^{\dagger}a-b^{\dagger}b$ with $a^{\dagger}$ and $b^{\dagger}$ being bosonic creation operators for two orthogonal quantum states. Also, our interaction Hamiltonian is similar to that by coupling an atomic ensemble to a light beam with the off-resonant Faraday interaction Bao et al. (2020); Duan et al. (2024), where $J_{z}$ is the sum of the total angular momentum of individual atoms and $\sigma_{z}$ denotes the Stokes operator associated with the distinction between the number operators of the photons polarized along orthogonal bases.
@@
Figure 1: Circuit model of our qubit-assisted metrology consisting of a two-step evolution interrupted by the parametric encoding and a single-shot measurement. Step $1$ and $2$ denote the free joint unitary evolution $U(t_{1})$ and its time reversal $U(t_{2})$, respectively, in between which a to-be-estimated parameter $\theta$ is encoded into the probe state via a unitary operation $R_{x}(\theta)$. The output state is determined by a projective measurement on either the full system or the ancillary qubit.
@@
The probe spin ensemble and the ancillary qubit are initially separable, i.e., the input state of the full system is a product state $\rho_{P}\otimes\rho_{A}$. The whole circuit in Fig. 1 can be described by
$U_{\theta}=U(t_{2})R_{x}(\theta)U(t_{1})=e^{-iHt_{2}}e^{-i\theta J_{x}}e^{-iHt_{1}},$ (2)
@@
where the two steps of evolution last $t_{1}$ and $t_{2}$, respectively. During Step $1$, the probe and ancillary system are entangled. Through the parametric encoding of a negligible evolution time, we obtain a phase parameter $\theta$ by a spin rotation $R_{x}(\theta)=\exp(-i\theta J_{x})$ of the probe system. Experimentally $\theta$ can be accumulated during the precessing about $z$-axis Meyer et al. (2001); Gross et al. (2010); Ockeloen et al. (2013) induced by a certain interaction between the probe and a to-be-measured system. Then the rotation about $x$-axis could be performed by a sequence of $R_{x}(\theta)=R_{y}(\pi/2)R_{z}(\theta)R_{y}(-\pi/2)$, where $R_{y}(\pm\pi/2)$ indicates a $\pi/2$ pulse about $y$-axis. The duration time of Step $2$ is determined by $t_{1}$ under the time-reversal strategy Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024). On the last stage of the circuit, a projective measurement about the full system is performed on the output state $\rho(\theta)\equiv U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}$, whose succuss probability can be used to infer the Fisher information about the estimated parameter. When the input state of the probe (spin ensemble) is an eigenstate of $U_{\theta}$, the maximum parameter information can be obtained only by performing a projective measurement on the ancillary qubit.
@@
It has been shown Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024) that the measurement precision can be certainly enhanced such that CFI is increased to be equivalent to QFI when the joint evolution operator $U(t_{2})$ becomes the time reversal of $U(t_{1})$, i.e.,
@@
It indicated that the full system traces back its entanglement generation trajectory and returns to the input state if no phase is encoded.
III Conditions for time reversal
@@
where $N=2j$ is the total spin number of the probe ensemble, $\mathcal{I}^{n}$ is the identity matrix with dimension $n$, $|j,m\rangle$ is the eigenstate of the operators $J_{z}$ with eigenvalue $m$, and $|g\rangle$ and $|e\rangle$ denote the ground and excited states of the ancillary spin, respectively. Using the full Hamiltonian (1), we have
@@
$\otimes(C_{e,m}|e\rangle\langle e|+C_{g,m}|g\rangle\langle g|),$
where $C_{e,m}=\exp\{-i[\omega_{A}+m(\omega_{P}+g)]T\}$ and $C_{g,m}=\exp\{i[\omega_{A}-m(\omega_{P}-g)]T\}$ with $T\equiv t_{1}+t_{2}$. It is found that when
@@
Eq. (4) becomes equivalent to Eq. (5) up to a global phase. Given the expression in Eq. (6), we have
@@
The parity of the probe-spin number $N$ determines whether the eigenvalue $m$ is an integer or a half-integer. When $N$ is even, $m$ is an integer and then $m\pm m^{\prime}$ are the same in parity. A sufficient solution for Eq. (7) is thus $e^{igT}=e^{i\omega_{P}T}=\pm 1$. Straightforwardly we have
@@
where $n_{j}$’s, $j=1,2,4$, are proper integers. For either even or odd, it might be hard to find the exact solution of $T$ if $g$, $\omega_{P}$, and $\omega_{A}$ are incommensurable.
@@
In Fig. 2, we demonstrate the normalized trace $\mathcal{F}(T)$ as a function of the evolution time $T$ exemplified with $N=4$ and $N=11$ under a fixed setting that $\omega_{P}=\omega_{A}=3g$. The parity of the probe-spin number $N$ determines the time when the normalized trace $\mathcal{F}(T)$ attains unit. When $N=4$ (the even case, see the blue solid line), it is found that the normalized trace attains the maximum value when the full evolution time $T$ is a multiple of $\pi/g$. For example, when $T=\pi/g$, Eq. (8) is valid for $n_{1}=1$, $n_{2}=1$, and $n_{3}=3$. When $N=11$ (the odd case, see the red dashed line), the normalized trace attains unit when $T$ is a multiple of $2\pi/g$. Particularly, $T=2\pi/g$ is attained by setting $n_{1}=2$, $n_{2}=2$, and $n_{4}=5$ in Eq. (9). Moreover in case of an odd $N$, one can find that the unitary operator becomes
@@
up to a global phase when $T$ is an odd multiple of $\pi/g$. It will give rise to a vanishing normalized trace $\mathcal{F}(T)=0$. And then the time-reversal strategy is broken. Therefore, choosing a probe ensemble with an even number of spins can significantly accelerate our metrology protocol.

### v2

@@ Abstract
We propose a quantum metrology protocol based on a two-step joint evolution of the probe system and an ancillary qubit and quantum measurement. With a proper initial state of the ancillary qubit and an optimized evolution time, the quantum Fisher information (QFI) about the phase parameter encoded in the probe system is found to be determined by the expectation value of the square of a phase generator, irrespective of the probe initial state. Consequently, even if the probe is prepared as a finite-temperature state, faraway from the so-called resource state, e.g., the squeezed spin state or the Greenberger-Horne-Zeilinger state in atomic systems, the QFI in our protocol can approach the Heisenberg scaling $N^{2}$ with respect to the probe size $N$. This quadratic scaling behavior shows robustness against the imperfections about the initial state of the ancillary qubit and the optimized evolution time of the whole system. Also it is not sensitive to the deviation in system parameters and the qubit decoherence. Using the time-reversal strategy and a single-shot projective measurement, the classical Fisher information (CFI) in our metrology protocol is saturated with its quantum counterpart. Our work thus provides an economical method to reach the Heisenberg limit in metrology precision with no input of entanglement or squeezing.
@@ I Introduction
Quantum metrology aims to use various quantum resources including entanglement and squeezing to go beyond the standard quantum limit (SQL), where the uncertainty or fluctuation decreases with $1/\sqrt{N}$, on performing measurements of observable parameters in a probe with $N$ elementary sub-units Sun et al. (2010); Ma et al. (2011); Genoni et al. (2012); Escher et al. (2012); Zhong et al. (2013); Giovannetti et al. (2004). Examples of these parameters are the phases of evolving quantum systems accumulated in multiple scenarios Paris (2009); Helstrom (1969); Yurke et al. (1986), such as gravitational wave detection Caves (1981), biological sensing Taylor and Bowen (2016); Mauranyapin et al. (2017), atomic clock Ludlow et al. (2015); Katori (2011), and magnetometry Jones et al. (2009). A conventional protocol for precision metrology reaching the Heisenberg limit (HL) Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020); Pezzè et al. (2018), where the uncertainty decreases with $1/N$, consists of a prepared probe system [in a maximally correlated state Bollinger et al. (1996), e.g., the Greenberger-Horne-Zeilinger (GHZ) state in atomic systems and the NOON state in photonic systems], a definite unitary transformation to encode the phase parameter $\theta$ into the probe state, a practical measurement, and a suitable data processing to produce an estimator $\tilde{\theta}(x)$ about the parameter $\theta$ from the outcome $x$. The estimation precision is quantified by the standard deviation $\delta\theta=\langle[\tilde{\theta}(x)-\theta]^{2}\rangle^{1/2}$, which is under the constraint of the Cramér-Rao inequality Helstrom (1969); Holevo (1982): $\delta\theta\geq 1/\sqrt{F_{c}}$ with $F_{c}$ the classical Fisher information (CFI). Maximizing $F_{c}$ over all possible measurements gives rise to the quantum Fisher information (QFI) $F_{Q}$ and hence the quantum Cramér-Rao bound (QCRB) $\delta\theta_{\rm min}=1/\sqrt{F_{Q}}$ Braunstein and Caves (1994); Braunstein et al. (1996); Luo (2003); Pezzé and Smerzi (2009); Kacprowicz et al. (2010) on the attainable sensitivity to estimate $\theta$.
@@
The metrology precision $1/\delta\theta_{\rm min}$ improves with the sub-unit number $N$ of the probe system. With the probe in separable states, we have QFI $\propto N$ for the standard quantum limit. With the probe in maximally entangled states, the QFI scaling can be promoted from linear to quadratic, attaining the Heisenberg scaling $F_{Q}\propto N^{2}$. Hence the metrology precision attains HL Leibfried et al. (2004); Mitchell et al. (2004); Boto et al. (2000); Gerry (2000). In atomic systems, so far an $18$-qubit GHZ state has been generated on a quantum processor Zhang et al. (2013); Song et al. (2019); Choi et al. (2014). Alternatively, the squeezed spin state can be used to reduce fluctuation to improve metrology precision Wineland et al. (1992); Kitagawa and Ueda (1993); Wineland et al. (1994); Goda et al. (2008); Sewell et al. (2012), which are typically generated by the one-axis twisting (OAT) and two-axis twisting (TAT) interactions Liu et al. (2011); Zhang et al. (2017). Collective OAT interaction $H_{\rm OAT}\propto J_{z}^{2}$ gives rise to a sub-HL noise-reduction $\propto 1/N^{2/3}$ for $N$ particles. This interaction is popular in nonlinear platforms, such as Bose-Einstein condensations (BEC) Gross et al. (2010); Riedel et al. (2010), trapped ions Bohnet et al. (2016); Lu et al. (2019), and superconducting qubits Song et al. (2019); Xu et al. (2020). For example, the metrological gain is about $\delta\theta/\delta\theta_{\rm SQL}\sim 0.4$ with $120$ atoms Gross et al. (2010). Under the TAT interaction $H_{\rm TAT}\propto J_{x}^{2}+J_{y}^{2}$, it is theoretically claimed that the squeezing degree can approach HL Zhang et al. (2017); Borregaard et al. (2017). Yet this interaction remains as a challenge in current platforms Helmerson and You (2001); Borregaard et al. (2017); Macrì et al. (2020). Quantum nondemolition measurement Pezzè et al. (2018); Hosten et al. (2016); Bao et al. (2020); Duan et al. (2024) can also generate spin squeezing of atomic ensemble. The metrological gain is found to be $\delta\theta/\delta\theta_{\rm SQL}\sim 0.8$ with $10^{5}$ laser-cooled atoms Sewell et al. (2012). In quantum optics, the archetype of a metrology experiment for phase estimation is the Mach-Zehnder interferometer. When beam splitters are replaced with optical parametric amplifiers, the resulting SU(1,1) interferometer has already proven useful in a HL fluctuation reduction only with coherent states Yurke et al. (1986); Li et al. (2014, 2016); Jing et al. (2011); Hudelist et al. (2014); Du et al. (2022). However, the photon loss eventually makes it even below SQL Marino et al. (2012).
@@
Recently, the coupling of the probe with an ancillary system is found to outperform SQL in the parametric estimation. In a protocol for measuring the light rotation with indefinite time direction Xia et al. (2024), an ancillary system serves as a quantum switch leading to a superposition of opposite rotations of the probe system. The rotation precision can thus be enhanced to HL even by using classical probe states. A precision of $12.9$ nrad on the light rotation measurement was experimentally realized with a $150$-order Laguerre-Gaussian beam. In the error-corrected quantum metrology Demkowicz-Dobrzański and Maccone (2014), the entanglement between the probe and an ancillary system can be used to reduce noise. Rapid and noiseless quantum operations jointly on the probe and the ancillary system between the encoding rounds empower Ref. Zhou et al. (2018) to perform error correction to suppress the damaging effects of the noise on the probe. By virtue of the time-reversal strategy Yurke et al. (1986); Pezzè et al. (2017); Agarwal and Davidovich (2022); Wang et al. (2024), projective measurement can be optimized to saturate QCRB for multiphase estimation Pezzè et al. (2017), given an encoded probe state. The interaction between a photonic system and an ancillary qubit can be manipulated by an external field to implement the unitary transformations along the forward- and reverse-time directions Luo and Yu (2023). The first generates a photonic entangled NOON state to improve QFI and the second allows CFI to saturate its quantum counterpart.
@@
In this work, we introduce a qubit-assisted metrology protocol to estimate the phase parameter $\theta$ imprinted to the probe system (a spin ensemble) during a spin rotation. It is based on a properly prepared ancillary qubit, the time-reversal strategy, and the projective measurement, but not on any resource state, e.g., the squeezed spin state or GHZ state Pezzè et al. (2018). The QFI of our protocol can be modified from the variance of the phase generator to the mean square of that with respect to the probe state. It therefore shows a Heisenberg scaling in terms of the total spin number even when the probe is prepared as a thermal state. The scaling behavior is not sensitive to the imprecise control over the initial state of the ancillary qubit, the optimized evolution time, and the system parameters. Qubit decoherence reduces QFI yet does not destroy the quadratic scaling. Moreover, CFI of our protocol is found to be coincident with QFI by the time-reversal strategy.
@@
The rest of this work is structured as follows. In Sec. II, we describe the circuit model of our metrology protocol with the ancillary qubit. In Sec. III, we investigate the conditions about the time-reversal strategy. Detailed derivations about the relevant unitary evolution operator and the phase generator are provided in Appendix A. In Sec. IV, we find that QFI in our protocol can be expressed as the mean square of an optimized phase generator with respect to the probe initial state. Detailed derivation about QFI under the deviations in probe-qubit coupling strength and probe eigenfrequency is provided in Appendix B. In Sec. V, we calculate CFI in our protocol as the extractable information from the probability distribution of the output state upon a single-shot projective measurement. In Sec. VI, we adapt our framework constructed in Secs. II, III, IV, and V to an alternative interaction between the probe system and an ancillary qubit. In Sec. VII, we discuss the behavior of QFI under the dephasing noise and then summarize the entire work.
II Metrology with ancillary qubit
@@
Consider a metrology model consisting of a giant spin probe (spin ensemble) and an ancillary spin-$1/2$. The full Hamiltonian including the free part $H_{0}$ and the interaction part $H_{I}$ reads ($\hbar\equiv 1$)
@@
The free Hamiltonian and the interaction Hamiltonian in Eq. (1) are feasible in various experiments. In quantum dots Gillard et al. (2022), $J_{z}$ and $\sigma_{z}$ describe the nuclear spins and the electron spin, respectively. In BEC, the Hamiltonian can be realized by two cavities containing two-component BECs coupled by an optical fiber Pyrkov and Byrnes (2013), where the operators $J_{z}$ and $\sigma_{z}$ correspond to the Schwinger boson operators $a^{\dagger}a-b^{\dagger}b$ in the relevant cavities with $a^{\dagger}$ and $b^{\dagger}$ being bosonic creation operators for two orthogonal quantum states. Our interaction Hamiltonian is also similar to that coupling an atomic ensemble to a light beam with the off-resonant Faraday interaction Bao et al. (2020); Duan et al. (2024), where $J_{z}$ is the sum of the total angular momentum of individual atoms and $\sigma_{z}$ denotes the Stokes operator associated with the distinction between the number operators of the photons polarized along orthogonal bases.
@@
Figure 1: Circuit model of our qubit-assisted metrology. Step $1$ and $2$ denote the free joint unitary evolution $U(t_{1})$ and its time reversal $U(t_{2})$, respectively, in between which a to-be-estimated parameter $\theta$ is encoded into the probe state via a unitary operation $R_{x}(\theta)$. The output state is determined by a projective measurement on the whole system or the ancillary qubit.
@@
The probe spin ensemble and the ancillary qubit are initially separable, i.e., the input state of the full system is a product state $\rho_{P}\otimes\rho_{A}$. It is popular in quantum metrology Chen et al. (2018a, b) and standard in open-quantum-system dynamics Szańkowski (2023). Practically, it is convenient to obtain an analytical expression for both QFI and CFI with the circuit model in Fig. 1. The entire evolution operator in the circuit can be described by
$U_{\theta}=U(t_{2})R_{x}(\theta)U(t_{1})=e^{-iHt_{2}}e^{-i\theta J_{x}}e^{-iHt_{1}}.$ (2)
@@
During Steps $1$ and $2$, the probe system and the ancillary qubit experience a joint time evolution lasting $t_{1}$ and $t_{2}$, respectively. Through the parametric encoding of a negligible evolution time (or the full Hamiltonian in between the two steps can be temporarily switched off), we obtain a phase $\theta$ of the probe system by a spin rotation $R_{x}(\theta)=\exp(-i\theta J_{x})$. Experimentally, $\theta$ can be accumulated during the precessing about the $z$ axis Meyer et al. (2001); Gross et al. (2010); Ockeloen et al. (2013) induced by a certain interaction between the probe and a to-be-measured system. Then the rotation about $x$ axis could be performed by a sequence of $R_{x}(\theta)=R_{y}(\pi/2)R_{z}(\theta)R_{y}(-\pi/2)$, where $R_{y}(\pm\pi/2)$ indicates the $\pi/2$ pulse about $y$ axis. The duration time $t_{2}$ of Step $2$ is determined by $t_{1}$ of Step $1$ under the time-reversal strategy Agarwal and Davidovich (2022); Wang et al. (2024). On the last stage of the circuit, a projective measurement about the full system is performed on the output state $\rho(\theta)\equiv U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}$, whose probability distribution can be used to infer the classical Fisher information about the estimated parameter. When the input state of the probe (spin ensemble) is an eigenstate of $U_{\theta}$, the entire parameter information can be obtained merely by performing a projective measurement on the ancillary qubit.
@@
It has been shown Yurke et al. (1986); Pezzè et al. (2017); Agarwal and Davidovich (2022); Luo and Yu (2023); Wang et al. (2024) that the measurement precision can be certainly enhanced such that CFI approaches QFI when the joint evolution operator $U(t_{2})$ becomes the time reversal of $U(t_{1})$, i.e.,
@@
It indicates that the full system traces back to the input (separable) state if no phase is encoded.
III Conditions on time reversal
@@
where $N=2j$ is the total spin number of the probe ensemble, $\mathcal{I}^{n}$ is the identity matrix of $n$ dimensions, $|j,m\rangle$ is the eigenstate of the operators $J_{z}$ with eigenvalue $m$, and $|g\rangle$ and $|e\rangle$ denote the ground and excited states of the ancillary spin, respectively. Using the full Hamiltonian (1), we have
@@
$\otimes\left(C_{e,m}|e\rangle\langle e|+C_{g,m}|g\rangle\langle g|\right),$
where $C_{e,m}=\exp\{-i[\omega_{A}+m(\omega_{P}+g)]T\}$ and $C_{g,m}=\exp\{i[\omega_{A}-m(\omega_{P}-g)]T\}$ with $T\equiv t_{1}+t_{2}$. When
@@
Eqs. (4) and (5) become equivalent to each other up to a global phase. Consequently, we have
@@
The parity of the probe-spin number $N$ determines whether the eigenvalue $m$ is an integer or a half-integer. When $N$ is even, $m$ is an integer and then $m\pm m^{\prime}$ are the same in parity. A sufficient solution for Eq. (7) is thus $e^{igT}=e^{i\omega_{P}T}=\pm 1$. Straightforwardly, we have
@@
where $n_{j}$’s, $j=1,2,4$, are proper integers. For either even or odd $N$, it might be hard to find the exact solution of $T$ if $g$, $\omega_{P}$, and $\omega_{A}$ are incommensurable.
@@
In Fig. 2, we demonstrate the normalized trace $\mathcal{F}(T)$ as a function of the evolution time $T$ exemplified with $N=4$ and $N=11$ under a fixed setting that $\omega_{P}=\omega_{A}=3g$. The parity of the probe-spin number $N$ determines the time when $\mathcal{F}(T)$ attains unit. For $N=4$ (the even case, see the blue solid line), it is found that the normalized trace attains the maximum value when the full evolution time $gT$ is a multiple of $\pi$. For example, when $gT=\pi$, Eq. (8) is valid for $n_{1}=1$, $n_{2}=1$, and $n_{3}=3$. For $N=11$ (the odd case, see the red dashed line), the normalized trace attains unit when $gT$ is a multiple of $2\pi$. Particularly, $gT=2\pi$ is attained by setting $n_{1}=2$, $n_{2}=2$, and $n_{4}=5$ in Eq. (9). Moreover, when both $N$ and $gT/\pi$ are odd integers, one can find that the unitary operator becomes
@@
up to a global phase. It will give rise to a vanishing normalized trace $\mathcal{F}(T)=0$. Then the time-reversal strategy is broken. Therefore, choosing a probe ensemble with an even number of spins can significantly accelerate our metrology protocol.